# SceneTracker: Long-term Scene Flow Estimation Network

### v1

@@ Abstract
Considering the complementarity of scene flow estimation in the spatial domain’s focusing capability and 3D object tracking in the temporal domain’s coherence, this study aims to address a comprehensive new task that can simultaneously capture fine-grained and long-term 3D motion in an online manner: long-term scene flow estimation (LSFE). We introduce SceneTracker, a novel learning-based LSFE network that adopts an iterative approach to approximate the optimal trajectory. Besides, it dynamically indexes and constructs appearance and depth correlation features simultaneously and employs the Transformer to explore and utilize long-range connections within and between trajectories. With detailed experiments, SceneTracker shows superior capabilities in handling 3D spatial occlusion and depth noise interference, highly tailored to the LSFE task’s needs. The code for SceneTracker is available at https://github.com/wwsource/SceneTracker.
@@ 1 Introduction
The existing work primarily focuses on scene flow estimation (SFE) and 3D object tracking (3D OT) tasks, which partially address the issue of online 3D motion capture. SFE aims to estimate the pixel-wise/point-wise 3D motion between consecutive stereo, RGB-D, or point cloud data, and it can capture fine-grained instantaneous 3D displacement. However, SFE is unable to capture the 3D trajectory over a period of time robustly. Due to the presence of occlusions, per-frame 3D displacements cannot form the trajectory through simple chaining. 3D OT aims to achieve long-term tracking of targets in point clouds and employs a 3D bounding box to mark the position and size of the target at each moment. It operates on the object level and provides limited information about 3D deformations within the object.
Considering the complementarity of scene flow estimation in the spatial domain’s focusing capability and 3D object tracking in the temporal domain’s coherence, this study aims to address a comprehensive new task that can simultaneously capture fine-grained and long-term 3D motion: long-term scene flow estimation (LSFE). LSFE seeks to estimate the query pixel/point’s 3D trajectory within a continuous stereo, RGB-D, or point cloud sequence online. Our task can be seen as an extension of SFE into a more extended temporal domain or an extension of 3D OT into the pixel/point-level spatial domain.
We propose SceneTracker, the first long-term scene flow estimation method, which takes a $T$-frame RGB-D video and the camera intrinsics as input, along with the starting 3D coordinates $x,y,z$ and starting time $t$ of an arbitrary 3D target to track, and produces a $T\times 3$ matrix as output, representing the positions in the camera coordinate system of the target across the given frames. SceneTracker adopts an iterative approach to approximate the optimal trajectory, overcoming the significant displacement challenges between frames. Besides, it dynamically indexes and constructs appearance and depth correlation features simultaneously, enhancing the ability to localize the target 3D position. Finally, it employs the Transformer to explore and utilize long-range connections within and between trajectories, further improving the accuracy of long-term scene flow estimation.
Compared to the scene flow-based chaining method and tracking any point-based depth indexing method, the proposed SceneTracker is highly tailored to the needs of the LSFE task. It can effectively address 3D spatial occlusion and depth noise interference, achieving median 3D error reductions of 39.0% and 85.9%, respectively.
@@
    To better capture the fine-grained and long-term 3D motion, a new task, long-term scene flow estimation, is studied.
@@
    A novel learning-based LSFE network, SceneTracker, is presented to estimate per-point 3D trajectory within an RGB-D video effectively.
@@
    Both qualitative and quantitative experiments are conducted to verify the validity of the proposed method.
@@ 2 Related Works
2.1 Optical Flow Estimation
Optical Flow Estimation (OFE) aims to estimate the pixel-wise 2D instantaneous motion between consecutive video frames. Many of its classic works have inspired approaches to solving the SFE task. FlowNet [2] is the first method using DNNs to estimate optical flow. After that, a series of methods, such as PWC-Net [17] and LiteFlowNet [4], adopt a pyramid structure to refine the estimated optical flow from coarse to fine. GMFlow [22] completely transforms the mainstream regression-based pipeline and reformulates the optical flow estimation problem as a global matching problem, identifying corresponding relationships by directly comparing feature similarities. RAFT [18] adopts an iterative method to continuously refine the optical flow results at the same high resolution and has achieved significant performance improvements. Since then, GMA [5] and SplatFlow [20] have optimized this single-resolution iterative method from a spatial or temporal perspective, further improving the accuracy.
@@ 2.2 Scene Flow Estimation
As a fundamental yet challenging problem in 3D computer vision, SFE has been extensively studied in recent years due to its numerous applications, including autonomous driving, human pose estimation, moving object segmentation, and many others. The existing methods generally rely on feature similarities as their underlying principle. FlowNet3D [9] operates directly on point clouds with PointNet++ [15]. CamLiFlow [8] refines the optical flow and scene flow from the two frames’ feature similarity tensor in a coarse-to-fine way with two branches. OpticalExp [23] uses optical expansion to recover the pixel’s motion in depth between frames. It can estimate the final depth value of the second frame by combining the first frame’s depth, indirectly estimating the scene flow per pixel. The existence of occlusion issues severely constrains the accuracy of SFE. To tackle the occlusion problem, RigidMask [24] and RAFT-3D [19] focus on introducing rigid motion priors explicitly and implicitly, respectively.
@@ 2.3 Tracking Any Point
Unlike OFE, which focuses on local instantaneous motion, the Tracking Any Point (TAP) task focuses more on the global motion trajectory in the image domain. PIPs [3] combines cost volumes and iterative inference with a deep temporal network, which jointly reasons about the location and appearance of visual entities across multiple timesteps, achieving excellent tracking results. TAP-Net [1] computes cost volumes for the first frame and each intermediate frame of the video independently, which will be directly used for regression of position and occlusion status. MFT [14] performs the chaining of optical flow and occlusion status by estimating the uncertainty. Cotracker [6] utilizes the strong correlation between multiple tracking points for joint tracking. PIPs++ [25] introduces 1D convolution and dynamic templates, improving the tracking survival rate over longer time spans. In contrast to the TAP task, LSF involves overcoming inherent noise in depth data and considering the 3D structure of objects and their motion patterns to accurately track trajectories in 3D space.
@@ 3 Proposed Method
In this section, the holistic architecture of our method is described in Section 3.1. Then, two key contributions of our method, including the Flow Iteration Module and the Transformer Cell network, are depicted in Section 3.2 and 3.3, respectively. Some details of the model during training and inference are provided in Section 3.4.
@@ 3.1 Overall Architecture of SceneTracker
For the first sliding window, the positions are initialized as the initial positions of the query points. For the other sliding windows, the first $S/2$ frames are initialized based on the estimated results of the last $S/2$ frames of the previous sliding window, and the last $S/2$ frames are all initialized based on the estimated result of the last frame of the previous sliding window. Notably, for the last sliding window, if its number of frames is less than $S$, we will temporarily duplicate the last frame data to make up $S$ frames. Now, taking any current sliding window $W_{c}$ as an example, we can obtain the initialized trajectories $P^{xyz}_{init}\in\mathbb{R}^{N\times T\times 3}$.
Feature Encoder and downsampling Our network is performing inference at resolution $\frac{H}{s}\times\frac{W}{s}$. Here, $s=8$ is a downsampling factor utilized for efficiency. First, we use a Feature Encoder network to extract image features $\overline{F}=\{\overline{F_{k}},...,\overline{F_{k+S-1}}\}$ from $\{\overline{I_{k}},...,\overline{I_{k+S-1}}\}$ in the current sliding window $W_{c}$, assuming that $W_{c}$’s first frame is the entire video’s $k$th frame. The Feature Encoder is a convolutional neural network comprising 8 residual blocks and 5 downsampling layers. The temporary features at 4 scales will generate a dense feature map at $1/s$ the input image resolution through scaling operations and convolutional layers.
@@
To match the generated $\overline{F}$ and $\overline{D}$, we downsample the initialized trajectory to generate the $P^{uvd}_{0}$:
$P_{0}^{uvd}=\left[\begin{array}[]{c}P_{0}^{u}\\ P_{0}^{v}\\ P_{0}^{d}\end{array}\right]=\left[\begin{array}[]{c}P_{init}^{u}/s\\ P_{init}^{v}/s\\ P_{init}^{d}\end{array}\right].$ (2)
Update of template feature and trajectory In the Flow Iteration Module, we iteratively update the query points’ template features and 3D trajectories. When processing the first sliding window’s first frame, we perform bilinear sampling on $\overline{F_{1}}$ using the $uv$ coordinates of the query points to obtain $\overline{F_{1}}$’s initial template feature. Then, we replicate this feature $S$ times along the temporal dimension to obtain the initial template feature $Q_{0}$ for all subsequent sliding windows. Each sliding window will have a consistent $Q_{0}$ and different $P_{0}^{uvd}$. After $n$ iterations of the same Transformer Predictor module, they will be updated to $Q_{n}$ and $P_{n}^{uvd}$.
Trajectory output We first upsample $P_{n}^{uvd}$ to generate $P^{uvd}$ to match the original input resolution:
$P^{uvd}=\left[\begin{array}[]{c}P^{u}\\ P^{v}\\ P^{d}\end{array}\right]=\left[\begin{array}[]{c}P_{n}^{u}*s\\ P_{n}^{v}*s\\ P_{n}^{d}\end{array}\right].$ (3)
Then, we combine the camera intrinsics $K$ to transform $P^{uvd}$ from the $uvd$ coordinate system to the camera coordinate system, resulting in $P^{xyz}$. The mathematical expressions of the conversion are as follows,
$P^{xyz}=\left[\begin{array}[]{c}P^{x}\\ P^{y}\\ P^{z}\end{array}\right]=\left[\begin{array}[]{c}P_{d}\cdot(P^{u}-c_{x})/f_{x}\\ P_{d}\cdot(P^{v}-c_{y})/f_{y}\\ P_{d}\end{array}\right].$ (4)
Finally, we chain all $P^{xyz}$ estimated from the sliding windows. The overlapping parts between adjacent windows take the results from the latter window.
@@ 3.2 Flow Iteration Module
Taking the FIM’s $n$th iteration as an example, the Transformer Predictor network requires inputs $Q_{n-1}$, $\overline{F}$, $P_{n-1}^{uvd}$, and $\overline{D}$ to construct the input for a Transformer Cell, which involves concatenating the apparent correlation feature $C_{n-1}^{uv}$, the depth correlation feature $C_{n-1}^{d}$, the apparent motion $O_{n-1}^{uv}$, and the depth motion $O_{n-1}^{d}$. The Transformer Cell outputs residuals, $\Delta Q_{n}$ and $\Delta P_{n}^{uvd}$, to update $Q_{n-1}$ and $P_{n-1}^{uvd}$ to $Q_{n}$ and $P_{n}^{uvd}$, serving as the output for the current iteration:
$P_{n}^{uvd}$ $=P_{n-1}^{uvd}+\Delta P_{n}^{uvd}$ (5)
$Q_{n}$ $=Q_{n-1}+\Delta Q_{n}$ (6)
@@
Depth correlation feature The network uses $P_{n-1}^{uvd}$ to lookup $\overline{D}$ to obtain $C_{n-1}^{d}$. Specifically, $C_{n-1}^{d}$ is obtained through bilinear sampling of $\overline{D}$’s inverse depth map with the grid size $(2\times r_{w}+1)\times(2\times r_{h}+1)$, calculating the residual with the depth prediction $P_{n-1}^{d}$’s inverse, and concatenating. Here, $r_{w}$ and $r_{h}$ are the local neighborhood radii in the $u$ and $v$ directions.
@@ 3.3 Transformer Cell Network
First, $x_{input}$ will be sinusoidally encoded separately in spatial and temporal dimensions to obtain $\eta_{space}$ and $\eta_{time}$. Then, we perform element-wise addition on $x_{input}$, $\eta_{space}$, and $\eta_{time}$ to obtain $x_{input}^{\eta}$. Afterward, $x_{input}^{\eta}$ will undergo $2\times M$ Transformer Blocks for feature enhancement, with Cross-time and Cross-space Transformer Blocks being executed alternately. The two types of Transformer Blocks divide $x_{input}^{\eta}$ into tokens based on trajectories and time points, respectively. After tokenization, both types of Transformer Blocks will sequentially undergo a normalization layer, multi-head self-attention layer, normalization layer, and MLP layer, accompanied by two skip connections. The model can effectively explore and utilize long-range connections within and between trajectories through the operations above.
@@ 3.4 Training and Inference
Here we set $\gamma=0.8$ and $\alpha=250$. $W_{max}$ is the last sliding window.

### v2

@@ Abstract
Considering the complementarity of scene flow estimation in the spatial domain’s focusing capability and 3D object tracking in the temporal domain’s coherence, this study aims to address a comprehensive new task that can simultaneously capture fine-grained and long-term 3D motion in an online manner: long-term scene flow estimation (LSFE). We introduce SceneTracker, a novel learning-based LSFE network that adopts an iterative approach to approximate the optimal trajectory. Besides, it dynamically indexes and constructs appearance and depth correlation features simultaneously and employs the Transformer to explore and utilize long-range connections within and between trajectories. With detailed experiments, SceneTracker shows superior capabilities in handling 3D spatial occlusion and depth noise interference, highly tailored to the LSFE task’s needs. Furthermore, we build the first real-world evaluation dataset, LSFDriving, to boost the LSFE field. The code for SceneTracker is available at https://github.com/wwsource/SceneTracker.
Index Terms:
Long-term Scene Flow Estimation, Scene Flow Estimation, 3D Object Tracking, Transformer
@@ 1 Introduction
The existing work primarily focuses on scene flow estimation (SFE) and 3D object tracking (3D OT) tasks, which partially address the issue of online 3D motion capture. SFE aims to estimate the pixel-wise/point-wise 3D motion between consecutive stereo, RGB-D, or point cloud data, and it can capture fine-grained instantaneous 3D displacement. However, SFE is unable to capture the 3D trajectory over a period of time robustly. Due to the presence of occlusions, per-frame 3D displacements cannot form the trajectory through simple chaining. 3D OT aims to achieve long-term tracking of targets and employs a 3D bounding box to mark the position and size of the target at each moment. It operates on the object level, requires prior modeling of the object [1, 2, 3], and provides limited information about 3D deformations within the object.
Considering the complementarity of SFE in the spatial domain’s focusing capability and 3D OT in the temporal domain’s coherence, this study aims to address a comprehensive new task that can simultaneously capture fine-grained and long-term 3D motion: long-term scene flow estimation (LSFE). LSFE seeks to estimate the query pixel/point’s 3D trajectory within a continuous stereo, RGB-D, or point cloud sequence online. Our task can be seen as an extension of SFE into a broader temporal domain or an extension of 3D OT into the pixel/point-level spatial domain.
We propose SceneTracker, the first LSFE method, which takes a $T$-frame RGB-D video and the camera intrinsics as input, along with the starting 3D coordinates $x,y,z$ and starting time $t$ of an arbitrary 3D target to track, and produces a $T\times 3$ matrix as output, representing the positions in the camera coordinate system of the target across the given frames. SceneTracker adopts an iterative approach to approximate the optimal trajectory, overcoming the significant displacement challenges between frames. Besides, it dynamically indexes and constructs appearance and depth correlation features simultaneously, enhancing the ability to localize the target 3D position. Finally, it employs the Transformer to explore and utilize long-range connections within and between trajectories, further improving the accuracy of LSFE.
We train and evaluate the proposed method on the large-scale synthetic dataset. Compared to the scene flow-based chaining method and tracking any point-based depth indexing method, the proposed SceneTracker is highly tailored to the needs of the LSFE task. It can effectively address 3D spatial occlusion and depth noise interference, achieving median 3D error reductions of 57.7% and 86.6%, respectively.
Moreover, we build a real-world evaluation dataset, LSFDriving. This dataset is geared towards the autonomous driving realm, providing 3D trajectory annotations for points on objects with diverse motion attributes and rigidities. Specifically, the tracked points are sampled from static backgrounds, moving rigid vehicles, and moving non-rigid pedestrians. Addressing the complexity of local non-linear motion in pedestrian joints and garments, we build a semi-automated annotation pipeline capable of efficiently and accurately annotating 3D trajectories.
@@
    To better capture the fine-grained and long-term 3D motion, a comprehensive new task, long-term scene flow estimation, is studied.
@@
    A novel learning-based LSFE network, SceneTracker, is presented. SceneTracker shows superior capabilities in handling 3D spatial occlusion and depth noise interference.
@@
    To boost the development of the LSFE field, we build the first real-world evaluation dataset, LSFDriving.
@@ 2.1 Scene Flow Estimation
The existing methods generally rely on feature similarities as their underlying principle. FlowNet3D [4] operates directly on point clouds with PointNet++ [5]. CamLiFlow [6] refines the optical flow and scene flow from the two frames’ feature similarity tensor in a coarse-to-fine way with two branches. OpticalExp [7] uses optical expansion to recover the pixel’s motion in depth between frames. It can estimate the final depth value of the second frame by combining the first frame’s depth, indirectly estimating the scene flow per pixel. The existence of occlusion issues severely constrains the accuracy of SFE. To tackle the occlusion problem, RigidMask [8] and RAFT-3D [9] focus on introducing rigid motion priors explicitly and implicitly, respectively.
@@ 2.2 Tracking Any Point
Unlike optical flow estimation methods [10, 11, 12, 13, 14, 15, 16], which focuses on local instantaneous motion, the tracking any point (TAP) task focuses more on the global motion trajectory in the image domain. TAP-Net [17] computes cost volumes for the first frame and every intermediate frame of the video independently, which will be directly used for regression of position and occlusion status. MFT [18] performs the chaining of optical flow and occlusion status by estimating the uncertainty. Cotracker [19] utilizes the strong correlation between multiple tracking points for joint tracking. PIPs++ [20] introduces 1D convolution and dynamic templates, improving the tracking survival rate over longer time spans in comparison to its previous version [21]. In contrast to the TAP task, LSFE involves overcoming inherent noise in depth data and considering the 3D structure of objects and their motion patterns to track trajectories in 3D space accurately.
@@ 3.1 Overall Architecture of SceneTracker
For the first sliding window, the positions are initialized as the initial positions of the query points. For the other sliding windows, the first $S/2$ frames are initialized based on the estimated results of the last $S/2$ frames of the previous sliding window, and the last $S/2$ frames are all initialized based on the estimated result of the last frame of the previous sliding window. Now, taking any current sliding window $W_{c}$ as an example, we can obtain the initialized trajectories $P^{xyz}_{init}\in\mathbb{R}^{N\times T\times 3}$.
Feature Encoder and downsampling Our network is performing inference at resolution $\frac{H}{s}\times\frac{W}{s}$. Here, $s=8$ is a downsampling factor utilized for efficiency. First, we use a Feature Encoder network to extract image features $\overline{F}=\{\overline{F_{k}},...,\overline{F_{k+S-1}}\}$ from $\{\overline{I_{k}},...,\overline{I_{k+S-1}}\}$ in the current sliding window $W_{c}$, assuming that $W_{c}$’s first frame is the entire video’s $k$th frame. The Feature Encoder is a convolutional neural network comprising 8 residual blocks and 5 downsampling layers.
@@
To match the generated $\overline{F}$ and $\overline{D}$, we downsample the initialized trajectory to generate the $P^{uvd}_{0}$.
Updates of template feature and trajectory In the Flow Iteration Module, we iteratively update the query points’ template features and 3D trajectories. When processing the first sliding window’s first frame, we perform bilinear sampling on $\overline{F_{1}}$ using the $uv$ coordinates of the query points to obtain $\overline{F_{1}}$’s initial template feature. Then, we replicate this feature $S$ times along the temporal dimension to obtain the initial template feature $Q_{0}$ for all subsequent sliding windows. Each sliding window will have a consistent $Q_{0}$ and different $P_{0}^{uvd}$. After $n$ iterations of the same Transformer Predictor module, they will be updated to $Q_{n}$ and $P_{n}^{uvd}$.
Trajectory output We first upsample $P_{n}^{uvd}$ to generate $P^{uvd}$ to match the original input resolution. Then, we combine the camera intrinsics $K$ to transform $P^{uvd}$ from the $uvd$ coordinate system to the camera coordinate system, resulting in $P^{xyz}$. Finally, we chain all $P^{xyz}$ estimated from the sliding windows. The overlapping parts between adjacent windows take the results from the latter window.
@@ 3.2 Flow Iteration Module
Figure 2: Example of producing LSFE annotations for pedestrian samples. We combine the left image and depth map sequence to form the RGB-D video. Once the left video’s 2D trajectory and disparity sequence are determined, we can utilize the binocular camera’s calibration parameters to convert them into the ground truth 3D trajectory.
@@
Taking the FIM’s $n$th iteration as an example, the Transformer Predictor network requires to use $Q_{n-1}$, $\overline{F}$, $P_{n-1}^{uvd}$, and $\overline{D}$ to construct the input for a Transformer Cell, which involves concatenating the apparent correlation feature $C_{n-1}^{uv}$, the depth correlation feature $C_{n-1}^{d}$, the apparent motion $O_{n-1}^{uv}$, and the depth motion $O_{n-1}^{d}$. The Transformer Cell outputs residuals, $\Delta Q_{n}$ and $\Delta P_{n}^{uvd}$, to update $Q_{n-1}$ and $P_{n-1}^{uvd}$ to $Q_{n}$ and $P_{n}^{uvd}$, serving as the output for the current iteration.
@@
Depth correlation feature The network uses $P_{n-1}^{uvd}$ to lookup $\overline{D}$ to obtain $C_{n-1}^{d}$. Specifically, $C_{n-1}^{d}$ is obtained through bi-linear sampling of $\overline{D}$’s inverse depth map and calculating the residual with the depth prediction $P_{n-1}^{d}$’s inverse.
@@ 3.3 Transformer Cell Network
First, $x_{input}$ will be sinusoidally encoded separately in temporal and spatial dimensions to obtain $\eta_{time}$ and $\eta_{space}$. Then, we perform element-wise addition on $x_{input}$, $\eta_{time}$, and $\eta_{space}$ to obtain $x_{input}^{\eta}$. Afterward, $x_{input}^{\eta}$ will undergo $2\times M$ Transformer Blocks for feature enhancement, with Cross-time and Cross-space Transformer Blocks being executed alternately. The two types of Transformer Blocks divide $x_{input}^{\eta}$ into tokens based on time points and trajectories, respectively. After tokenization, both types of Transformer Blocks will sequentially undergo a normalization layer, multi-head self-attention layer, normalization layer, and MLP layer, accompanied by two skip connections. The model can effectively explore and utilize long-range connections within and between trajectories through the operations above.
@@ 3.4 Training and Inference
Here we set $\gamma=0.8$ and $\alpha=250$. $W_{max}$ is the final sliding window.
4 Proposed Dataset
Generally, sparse annotations are only available for rigid bodies across adjacent frames in real-world datasets [25, 26] of SFE. Differing from these datasets, the proposed LSFE dataset requires additional annotations of trajectories that span across multiple frames and are not limited to rigid bodies. Given a temporal segment of autonomous driving data, we aim to construct a $T$-frame RGB-D video and the 3D trajectory of interested points in the first frame. Specifically, the interested points are sampled from static backgrounds, moving rigid vehicles, and moving non-rigid pedestrians.
4.1 Annotations on Backgrounds
First, we utilize the camera’s intrinsic and extrinsic parameters to extract the first frame’s LiDAR points, which can be projected correctly into the image. Then, we use the bounding boxes from 2D object detection to exclude all foreground LiDAR points. Here, we have obtained the first frame’s background LiDAR points. Taking one LiDAR point $X\in\mathbb{R}^{3}$ as an example, we project it onto the remaining $T-1$ frames using vehicle poses. Formally, the projected point at the time $t$ is:
$X_{t}=W_{t}^{-1}\cdot W_{1}\cdot X$ (6)
Here, $W_{t}$ represents the transformation from the ego vehicle to the world coordinate system at the time $t$.
4.2 Annotations on Vehicles
In contrast to the background, the vehicle exhibits its own distinctive motion. To account for this, we follow [27] to introduce 3D bounding boxes from 3D object tracking to provide the transformation $B_{t}$ from the world to the box coordinate system at the time $t$. We use the 3D bounding boxes to filter out all vehicle LiDAR points. Taking one LiDAR point $X\in\mathbb{R}^{3}$ as an example, the projected point at the time $t$ is:
$X_{t}=W_{t}^{-1}\cdot B_{t}^{-1}\cdot B_{1}\cdot W_{1}\cdot X$ (7)
4.3 Annotations on Pedestrians
The intricacy and non-rigidity of pedestrian movement determine their annotation difficulties, as evidenced by its absence in existing SFE datasets. We employ binocular video to address this challenge indirectly. Firstly, we prepare a segment of T-frames of rectified binocular video.
Then, we employ a semi-automated annotation pipeline to efficiently and accurately label the 2D trajectories of interested points in the left and right view videos.
Labeling the interested point We develop a bespoke annotation software and label the 2D coordinates $x_{1}^{l}$ of the interested points within the left image of the first frame.
Comuting the coarse left trajectory We utilize Cotracker [19] to compute $x_{1}^{l}$’s coarse 2D trajectory in the left-view video. Maintaining the first frame’s coordinates, we obtain the trajectory $S^{l}=\{x_{1}^{l},...,x_{T}^{l}\}$.
Comuting the coarse right trajectory We employ LEAStereo [28] to compute the disparities $\{c_{1},...,c_{T}\}$ of the interested point frame by frame. Combining the disparities and $S^{l}$, we derive the coarse right trajectory $S^{l}=\{x_{1}^{l},...,x_{T}^{l}\}$.
Manual refinement phase The left and right coarse trajectories will be displayed in the annotation software, where the human annotators will correct any low-quality annotations.
Finally, we combine the refined left trajectory with the disparity sequence to construct the 3D trajectory. The depth at frame $t$ is:
$d_{t}=f_{x}\cdot b/c_{t}$ (8)
Here $b$ is the binocular baseline length. Figure 2 illustrates an example of producing LSFE annotations for pedestrian samples.
Figure 3: Visualizations of our SceneTracker’s estimation results on the LSFOdyssey test dataset. The predicted 3D trajectories are depicted in blue on their corresponding point clouds.